x squared minus 2x equals 8
Answer by Cromlix(4381)

Hi there,
x^2 - 2x = 8
x^2 - 2x - 8 = 0
(x - 4)(x + 2) = 0
x - 4 = 0
x = 4
x + 2 = 0
x = -2
Hope this helps :-)